# A neural network-based approach to hybrid systems identification for control

### v1

@@ 4 The OCP in (1) is a data-based MPCC
We now derive an equivalent formulation for the OCP in (1), which will shed further light on the type of control problem we are dealing with. We will then present in §4.2 available results that are applicable to MPCCs models with the particular structure we obtain, before applying those results to our problem in §5.
@@ 4.1 Mathematical derivation
Suppose that we have trained $\mathcal{N}_{\theta}$ in (3) to obtain the structure of the OptNet layer as in (4) with $Q\in\mathbb{S}^{s}_{\succ 0}$ and imposing $c=0$. The following KKT conditions are then necessary and sufficient to characterize the optimal solution $z^{\star}$:
@@
where $\lambda\in\mathbb{R}^{l}_{\geq 0}$ represents the vector of Lagrange multipliers associated for the linear constraints. By recalling that $y=\textrm{col}(x,u)$, from the KKT system above we obtain a so-called LC model of the system dynamics [33] associated with the NN $\mathcal{N}_{\theta}$, with architecture as in Fig. 1:
@@
Note that all of the terms in (6) except $(x,u)$ and $\lambda$ are determined according to some offline training process characterizing the NN $\mathcal{N}_{\theta}$ in (3). Hence the terminology “data-based representation”.
@@
where the matrices $A$, $B_{u}$, $B_{w}$, $E_{w}$, $E_{x}$, $E_{u}$, and vectors $d$ and $e$ can be obtained via straightforward rearrangement of terms and definition of the complementarity variable $w\coloneqq\lambda$. Specifically, $B_{w}\coloneqq-WQ^{-1}F^{\top}$, $d\coloneqq-WQ^{-1}p$, $E_{w}\coloneqq FQ^{-1}F^{\top}$, $e\coloneqq FQ^{-1}p+h$. The matrices $A$ and $B_{u}$ (respectively, $E_{x}$ and $E_{u}$) are obtained by partitioning the matrix $-WQ^{-1}R$ (resp., $FQ^{-1}R-G$) with $n+m$ columns into two matrices with $n$ and $m$ columns, respectively. The notation is then analogous to (1), and we denote by $\boldsymbol{w}\coloneqq\textrm{col}((w_{k})_{t\in\mathcal{T}})$ the trajectory of complementarity variables corresponding to $\boldsymbol{u}$ and $\boldsymbol{x}$. A distinct feature of the LC dynamics are the complementarity constraints inherited from (6), which therefore turns the OCP (7), equivalent to (1) under our choice of $\mathcal{N}_{\theta}$ in (3), into a data-based MPCC.
@@
Unlike [44], where an LC model must be inferred from data directly, we obtain it as a consequence of choosing to train a NN $\mathcal{N}_{\theta}$ with architecture as in Fig. 1. For this reason, we do not require strict complementarity of the resulting LC model to recover differentiability w.r.t. the main parameters, as postulated instead in [44, 18].
@@ 4.2 Prior results on a class of MPCCs
In general, MPCCs amount to nonlinear, nonconvex optimization problems that can be very challenging to solve [56]. Indeed, for such problems the standard constraint qualifications for nonlinear programming (NLP) (e.g., the classical Mangasarian-Fromovitz one [8, §5.5.4]) typically fail to hold at any feasible point [14].
On the other hand, it has been recently proven in [38] that if the MPCC has a specific structure then one is able to establish strong stationarity conditions characterizing a (local) optimal solution. In particular, by referring to the data-based MPCC in (7) the following three requirements on the LC model are sufficient to recover the strong stationarity conditions derived in [38, Th. 1]:
@@
We consider an elementwise decomposition of the complementarity problem $0\leq E_{w}w+E_{x}x+E_{u}u+e\perp w\geq 0$ for simplicity, although a generalization of Condition 2 allowing for a block-diagonal decomposition is also possible – see [38, Ass. 2].
@@
([38, Th. 1]) Let $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ be feasible for the MPCC in (7) with an LC model satisfying Conditions 13. Then, $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ is locally optimal if and only if the standard KKT conditions for (7) admit a primal-dual solution pair.
The statement in Lemma 1 enables one to seek a local solution to the optimal control MPCC (7) through the solution of a classical KKT system, i.e., as an NLP. This is typically computationally advantageous with significantly better scaling features compared to more traditional mixed-integer approaches to solving OCPs based on generic PWA or hybrid models111 The MPCC in (7) is an OCP with hybrid dynamics, and consequently can be shown to be NP-hard. We therefore do not expect globally optimal solutions to be efficiently computable in general [17], regardless of the method employed. [38, 32].
Thus, if we are able to train our NN $\mathcal{N}_{\theta}$ in (3) so that Conditions 13 are satisfied, then $\mathcal{N}_{\theta}$ would also bring major computational advantages when used as part of an OCP.
@@ 5 Main results
Note that the above can always be satisfied for a given PWA system, e.g., by choosing any index $i\in\{1,\ldots,n_{r}^{\alpha}\}$ and $j\in\{1,\ldots,n_{r}^{\beta}\}$ and by setting
@@
Figure 2: Proposed NN-based hybrid system identification for control approach. After choosing hyperparameters $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ and diagonal $H\in\mathbb{S}^{2n}_{\succ 0}$, one directly obtains elements $Q$, $F$, $W$, and $c$ to be used, along with the available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, to train the NN $\mathcal{N}_{\theta}$ in (3) and (4). Plugging all the elements in (7) then enables to find some locally optimal triplet $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ by solving the KKT system associated to the underlying OCP problem.
By taking the NN-based perspective proposed in this work, however, one does not need to compute all of the aforementioned elements explicitly. As will be made clear in our main result, once the number of regions determining each partition is fixed, i.e., $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ (which will therefore coincide with the hyperparameters of our data-driven approach, along with matrices $Q_{\alpha}$, $Q_{\beta}$), we will obtain matrices and vectors $\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{n_{r}^{\alpha}}$, $\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{n_{r}^{\beta}}$, $(A_{\psi},B_{\psi},c_{\psi})$ and $(A_{\phi},B_{\phi},c_{\phi})$, as weights of the NN $\mathcal{N}_{\theta}$ in (3). In particular, we can claim the following:
@@
Let $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ be fixed. The NN $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ in (3), consisting of an OptNet layer and a buffer one in cascade, can be trained to produce a hybrid system model satisfying Conditions 13.
@@
To start we note that, for given $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ and any state-input pair $(x,u)\in\Omega$, (9b) and (9c) can be consolidated to obtain the following separable QP:
@@
where, for simplicity, we have imposed $\mathrm{diag}(Q_{\alpha},Q_{\beta})=H\in\mathbb{S}^{2n}_{\succ 0}$. Then, by introducing the decision variable $\xi\coloneqq\textrm{col}(\alpha,\beta)$ and after suitably redefining all matrices/vectors, the QP above admits the following compact reformulation:
@@
Theorem 1 asserts that, under a careful choice of some of the weights characterizing the NN $\mathcal{N}_{\theta}$ in (3), the latter can be trained by immediately means of samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ so that a data-based LC model meeting Conditions 13 is actually produced. In particular, our data-driven approach reduces the number of parameters that must be set during the training phase of $\mathcal{N}_{\theta}$ to $\{R,p,G,h\}$ – see Fig. 2 for a schematic representation of the proposed NN-based hybrid system identification for control mechanism.
@@
Corollary 1 requires that if our NN model $\mathcal{N}_{\theta}$ is to be included in the OCP (7), its parameters must also satisfy the strict inequalities in (10) in order for the local optimality results of Lemma 1 to apply.
However, these inequalities can not be enforced uniformly over $\Omega$ during the training phase, i.e., strict inequalities are guaranteed to hold true for the considered (state-input)-successor state samples only, $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, and not for all $(x,u)\in\Omega$. The relations in (11) provide us with a possible means of resolution by setting $(A_{\psi},B_{\psi},c_{\psi})$ and $(A_{\phi},B_{\phi},c_{\phi})$ to meet (10). Note that imposing equalities in (11) can be done either before or after the training process of $\mathcal{N}_{\theta}$.
@@
In [63] a strong stationarity condition is identified as necessary and sufficient to characterize local optima of a program with a rectified linear unit (ReLU) entering both in the cost and constraints, and the corresponding LC-based, lifted reformulation yielding an MPCC. In contrast, we give an explicit expression of a NN $\mathcal{N}_{\theta}$ for which local stationarity conditions, coinciding with the standard KKT system, are known to hold true for the MPCC obtained once employed the NN $\mathcal{N}_{\theta}$ as a data-based hybrid system model in the OCP (1).
@@ 6 Numerical experiments
We now present several examples in which the proposed NN $\mathcal{N}_{\theta}$ in (3)–(4) is adopted to learn continuous PWA models from dataset produced by PWA and nonlinear dynamics. In particular, in §6.1, we compare our approach with piecewise-affine regression and classification (PARC) algorithm [6]. In §6.2 we analyze how our technique is sensible to hyperparameters tuning. Finally, in §6.3 we make use of our NN-based model to solve OCPs with sensitivity-based solvers.
To do so, we focus on the structure reported in (9) directly, since Theorem 1 shows that the two representations are indeed equivalent, with the possibility to recover (3)–(4) from (9) and viceversa. Specifically, learning the latter amounts to determine the elements $\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{n_{r}^{\alpha}}$, $\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{n_{r}^{\beta}}$, $(A_{\psi},B_{\psi},c_{\psi})$, and $(A_{\phi},B_{\phi},c_{\phi})$, with hyperparameters $n_{r}^{\alpha}$ and $n_{r}^{\beta}$. With this regard, we note that also the choice of the Hessian matrix in (9b) and (9c), i.e., $Q_{\alpha}$ and $Q_{\beta}$, is arbitrary. Although they could also be included in the learning process, in the following we start by setting $Q_{\alpha}=Q_{\beta}=I$, and $n_{r}^{\alpha}=n_{r}^{\beta}=7$ unless otherwise stated. It is also interesting to note that, in this case, mappings $\alpha^{\star}(x,u)$ and $\beta^{\star}(x,u)$ in (9b) and (9c), respectively, recover the structure proposed in [50].
We implement the proposed approach in Python, utilizing JAX [10], JaxOPT [9] and Scipy [60]. The models have been trained to minimize the sum of the mean squared error and a regularization term penalizing the Euclidean norm of the parameter vector $\theta$. The overall training problem is therefore given by:
@@
where $\alpha_{\theta}(x,u)$ and $\beta_{\theta}(x,u)$ reminds the explicit dependence from $\theta=\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{7}\cup\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{7}\cup\{(A_{\psi},B_{\psi},c_{\psi}),(A_{\phi},B_{\phi},c_{\phi})\}$ of (9b) and (9c), respectively. The training problem was solved using the SLSQP algorithm [45] with $\lambda=0.01$. As internal QP solver, we relied on the $\mathtt{BoxCDQP}$ solver from JaxOPT. We then evaluate the approximation quality of the models obtained on the basis of open-loop predictions, obtained by simulating the trained $\mathcal{N}_{\theta}$ when excited with an unseen input sequence for 1000 time steps, and employing standard metrics [11, 13, 53, 51] such as best fit ratio (BFR):
$\mathrm{BFR}=\frac{\|\hat{X}^{+}-X^{+}\|_{F}}{\|X^{+}-\bar{X}^{+}\|_{F}},$
and the root mean square (RMS) error, defined as:
$\mathrm{RMS}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|x^{+,(i)}-\hat{x}^{+,(i)}\|_{2}^{2}}~{}.$
Referring to BFR, all matrices have dimension $n\times N$ and, specifically, $X^{+}$ stacks the true data, $\bar{X}^{+}$ is the componentwise average vector of $x$, while $\hat{X}^{+}$ is the predicted value of $\hat{X}^{+}$ obtained in open-loop prediction.
Unless otherwise stated, all tests have been conducted by making using 5000 normalized samples. We include white noise $N(0,0.01)$ to the sequence of state measurements to simulate the presence of measurement disturbances.

### v2

@@ 4 The OCP in (1) is a data-based MPCC
We derive next an equivalent formulation for the OCP in (1). Successively, we will then present in §4.2 available results that are applicable to MPCCs models with the particular structure we obtain, before applying those results to our problem in §5.
@@ 4.1 Mathematical derivation
Suppose that we have trained $\mathcal{N}_{\theta}$ in (3) to obtain the structure of the OptNet layer as in (4) with $Q\in\mathbb{S}^{s}_{\succ 0}$, and with $c=0$. The following KKT system is then necessary and sufficient to characterize the optimal solution $z^{\star}$:
@@
where $\lambda\in\mathbb{R}^{l}_{\geq 0}$ represents the vector of Lagrange multipliers associated to the linear constraints. By recalling that $y=\textrm{col}(x,u)$, from the KKT system above we obtain a so-called LC model of the system dynamics [29] associated with the NN $\mathcal{N}_{\theta}$, with architecture as in Fig. 1:
@@
Note that all of the terms in (6) except $(x,u)$ and $\lambda$ are determined through a training procedure for $\mathcal{N}_{\theta}$ in (3). Hence the terminology “data-based representation”.
@@
where the matrices $A$, $B_{u}$, $B_{w}$, $E_{w}$, $E_{x}$, $E_{u}$, and vectors $d$ and $e$ can be obtained by rearranging the terms and definition of the complementarity variable $w\coloneqq\lambda$. Specifically, $B_{w}\coloneqq-WQ^{-1}F^{\top}$, $d\coloneqq-WQ^{-1}p$, $E_{w}\coloneqq FQ^{-1}F^{\top}$, $e\coloneqq FQ^{-1}p+h$. The matrices $A$ and $B_{u}$ (respectively, $E_{x}$ and $E_{u}$) follow by partitioning $-WQ^{-1}R$ (resp., $FQ^{-1}R-G$) with $n+m$ columns into two matrices with $n$ and $m$ columns, respectively. The notation is then analogous to (1), and $\boldsymbol{w}\coloneqq\textrm{col}((w_{k})_{t\in\mathcal{T}})$ denotes the trajectory of complementarity variables corresponding to $\boldsymbol{u}$ and $\boldsymbol{x}$. A distinct feature of the LC dynamics are the complementarity constraints inherited from (6), which hence turns the OCP (7), equivalent to (1) under our choice of $\mathcal{N}_{\theta}$ in (3), into a data-based MPCC.
@@
Unlike [39], where an LC model shall be inferred from data directly, we obtain it as a consequence of choosing to train a NN $\mathcal{N}_{\theta}$ with architecture as in Fig. 1. For this reason, we do not require strict complementarity of the resulting LC model to recover differentiability w.r.t. the main parameters, as postulated instead in [39, 16].
@@ 4.2 Prior results on a class of MPCCs
In general, MPCCs amount to nonlinear, nonconvex optimization problems that can be very challenging to solve [50]. Indeed, for such problems the standard constraint qualifications for nonlinear programming (NLP) (e.g., the classical Mangasarian-Fromovitz one [7, §5.5.4]) typically fail to hold at any feasible point [12].
On the other hand, it has been recently proven in [33] that if the MPCC has a specific structure then one is able to establish strong stationarity conditions characterizing a (local) optimal solution. In particular, by referring to the data-based MPCC in (7) the following three requirements on the LC model are sufficient to recover the strong stationarity conditions derived in [33, Th. 1]:
@@
We consider an elementwise decomposition of the complementarity problem $0\leq E_{w}w+E_{x}x+E_{u}u+e\perp w\geq 0$ for simplicity, although a generalization of Condition 2 allowing for a block-diagonal decomposition is also possible – see [33, Ass. 2].
@@
([33, Th. 1]) Let $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ be feasible for the MPCC in (7) with an LC model satisfying Conditions 13. Then, $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ is locally optimal if and only if the standard KKT conditions for (7) admit a primal-dual solution pair.
The statement in Lemma 1 enables one to seek a local solution to the optimal control MPCC (7) through the solution of a classical KKT system, i.e., as an NLP. This is typically computationally advantageous with significantly better scaling features compared to more traditional mixed-integer approaches to solving OCPs based on generic PWA or hybrid models111 For NP-hard problems as (7) one should not expect globally optimal solutions to be efficiently computable in general [15], regardless of the method employed. [33, 28].
Then, if we manage to train our NN $\mathcal{N}_{\theta}$ in (3) so that Conditions 13 are satisfied, $\mathcal{N}_{\theta}$ would also bring major computational advantages when used as part of an OCP.
@@ 5 Main results
As suggested in [33], note that the above can always be satisfied for a given PWA system, e.g., by choosing any index $i\in\{1,\ldots,n_{r}^{\alpha}\}$ and $j\in\{1,\ldots,n_{r}^{\beta}\}$ and by setting
@@
Figure 2: Proposed NN-based hybrid system identification for control approach. After choosing hyperparameters $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ and diagonal $H\in\mathbb{S}^{2n}_{\succ 0}$, one directly obtains elements $Q$, $F$, $W$, and $c$ to be used, along with the available data $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, to train the NN $\mathcal{N}_{\theta}$ in (3) and (4) – see Theorem 1. Plugging all the elements in (7) then enables to find some locally optimal triplet $\left(\boldsymbol{x}^{\star},\boldsymbol{u}^{\star},\boldsymbol{w}^{\star}\right)$ by solving the KKT system associated to the underlying OCP problem.
By taking the NN-based perspective of in this work, however, one does not need to compute all of the aforementioned elements explicitly. As will be made clear in our main result, once the number of regions determining each partition is fixed, i.e., $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ (which will hence coincide with the hyperparameters of our data-driven approach, along with $Q_{\alpha}$, $Q_{\beta}$), we will obtain matrices and vectors $\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{n_{r}^{\alpha}}$, $\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{n_{r}^{\beta}}$, $(A_{\psi},B_{\psi},c_{\psi})$ and $(A_{\phi},B_{\phi},c_{\phi})$, as weights of $\mathcal{N}_{\theta}$ in (3). Specifically, we can claim the following:
@@
Let $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ be fixed. The NN $\mathcal{N}_{\theta}:\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ in (3) can be trained to produce a hybrid system model satisfying Conditions 13.
@@
To start we note that, for given $n_{r}^{\alpha}$, $n_{r}^{\beta}\geq 1$ and any state-input pair $(x,u)\in\Omega$, (9b) and (9c) can be lumped together to obtain the following separable QP:
@@
where, for simplicity, we have imposed $\mathrm{diag}(Q_{\alpha},Q_{\beta})=H\in\mathbb{S}^{2n}_{\succ 0}$. Then, by introducing the decision variable $\xi\coloneqq\textrm{col}(\alpha,\beta)$ and after suitably redefining all matrices/vectors, the QP above admits a compact form as:
@@
Under a careful choice of some of the weights characterizing the NN $\mathcal{N}_{\theta}$ in (3), the latter can then be trained by means of samples $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$ so that a data-based LC model meeting Conditions 13 is actually produced. In particular, our data-driven approach reduces the number of parameters that must be set during the training phase of $\mathcal{N}_{\theta}$ to $\{R,p,G,h\}$ – see Fig. 2 for a schematic representation. Alternatively, one may also decide to train the structure in (9) directly, and then successively recover the weights in (3)–(4). Specifically, learning the model in (9) amounts to determine the elements $\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{n_{r}^{\alpha}}$, $\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{n_{r}^{\beta}}$, $(A_{\psi},B_{\psi},c_{\psi})$, and $(A_{\phi},B_{\phi},c_{\phi})$, with hyperparameters $n_{r}^{\alpha}$ and $n_{r}^{\beta}$. To this end, note that also the choice of the Hessian matrix in (9b) and (9c), i.e., $Q_{\alpha}$ and $Q_{\beta}$, is arbitrary.
@@
Corollary 1 requires that if our NN model $\mathcal{N}_{\theta}$ is to be included in the OCP (7), its parameters must also satisfy the strict inequalities in (10) in order for the local optimality results of Lemma 1 to apply. However, as common to any type of NN, these inequalities can not be enforced uniformly over $\Omega$ during the training phase, i.e., strict inequalities are guaranteed to hold true for the considered (state-input)-successor state samples only, $\{(x^{(i)},u^{(i)},x^{+,{(i)}})\}_{i=1}^{N}$, and not for all $(x,u)\in\Omega$. Therefore, some a-posteriori verification procedure shall be applied to make sure they are satisfied for all $(x,u)\in\Omega$. Conversely, the relations in (11) provide us with a possible means of resolution by setting $(A_{\psi},B_{\psi},c_{\psi})$ and $(A_{\phi},B_{\phi},c_{\phi})$ to meet (10). In particular, imposing equalities in (11) can be done either before or after the training process of $\mathcal{N}_{\theta}$. Also in this case, however, applying some a-posteriori verification procedure is inevitable.
@@
In contrast to [55], we give an explicit expression of a NN $\mathcal{N}_{\theta}$ for which local stationarity conditions, coinciding with the standard KKT system, are known to hold true for the MPCC obtained once employed the NN $\mathcal{N}_{\theta}$ as a data-based hybrid system model in the OCP (1).
@@ 6 Numerical experiments
We now verify the performance of $\mathcal{N}_{\theta}$ in (3)–(4) to learning continuous PWA models from dataset produced by PWA and nonlinear dynamics. All simulations are run on a laptop equipped with a AMD R7 8840U processor.
To this end, we focus on the structure reported in (9) directly, since Theorem 1 establishes that the two representations are equivalent, and we set $Q_{\alpha}=Q_{\beta}=I$, and $n_{r}^{\alpha}=n_{r}^{\beta}=7$ unless otherwise stated. We implement our method in Python, utilizing JAX [9], JaxOPT [8] and Scipy [52]. The models have been trained to solve the following standard regularized training problem:
@@
where $\alpha_{\theta}(x,u)$ and $\beta_{\theta}(x,u)$ reminds the explicit dependence from $\theta=\{(A_{\alpha,i},B_{\alpha,i},c_{\alpha,i})\}_{i=1}^{7}\cup\{(A_{\beta,j},B_{\beta,j},c_{\beta,j})\}_{j=1}^{7}\cup\{(A_{\psi},B_{\psi},c_{\psi}),(A_{\phi},B_{\phi},c_{\phi})\}$ of (9b) and (9c), respectively. The training problem was solved using the SLSQP algorithm [40] with $\lambda=0.01$, and $\mathtt{BoxCDQP}$ from JaxOPT as internal QP solver.
We then evaluate the approximation quality of the models obtained on the basis of open-loop predictions, produced by exciting the trained $\mathcal{N}_{\theta}$ with an unseen input sequence for 1000 time steps, and employing standard metrics [10, 11, 47, 45] such as best fit ratio (BFR), defined as $\mathrm{BFR}=\|\hat{X}^{+}-X^{+}\|_{F}/\|X^{+}-\bar{X}^{+}\|_{F},$ and the root mean square (RMS) error, $\mathrm{RMS}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|x^{+,(i)}-\hat{x}^{+,(i)}\|_{2}^{2}}~{}.$ Referring to BFR, all matrices have dimension $n\times N$, with $X^{+}$ stacking the true data, $\bar{X}^{+}$ is the componentwise average vector of $x$, while $\hat{X}^{+}$ is the predicted value of $\hat{X}^{+}$ obtained in open-loop prediction. For system identification purposes, a high (respectively, low) value for BFR (RMS) is desirable.
Unless otherwise stated, all tests have been conducted by using 5000 normalized samples, adding white noise $N(0,0.01)$ to the sequence of state measurements.